# Random ReLU Neural Networks as Non-Gaussian Processes

### v1

@@ 1 Introduction
A shallow neural network is a function of the form
@@ 2 Generalized Stochastic Processes
The mathematical framework used in this paper is based on the theory of generalized stochastic processes (Itô, 1954; Gelfand, 1955; Gelfand and Vilenkin, 1964; Itô, 1984) as opposed to the more common “time-series” approach to studying stochastic processes. In this section, we present the relevant background on generalized stochastic processes. While this theory relies on some rather heavy concepts from functional analysis, it allows for elegant arguments to investigate the properties of the stochastic processes realized by the random neural networks in Eqs. 2 and 3.
@@ 4 Random ReLU Neural Networks as Stochastic Processes
where $\widetilde{{\bm{w}}}={\bm{w}}/\lVert{\bm{w}}\rVert_{2}$ and $\widetilde{b}=b/\lVert{\bm{w}}\rVert_{2}$. Therefore, the space of functions representable by shallow ReLU neural networks with input weights constrained to be unit norm is the same as the space of functions representable by shallow ReLU neural networks without constraints on the weights (Parhi and Nowak, 2023b). To that end, we focus on neurons of the form $\operatorname{ReLU}({\bm{w}}^{\mathsf{T}}{\bm{x}}-b)$ with $({\bm{w}},b)\in{\mathbb{S}^{d-1}\times\mathbb{R}}$.
@@
The equality in Eq. 18 is understood in $\mathcal{M}^{\mathsf{e}}({\mathbb{S}^{d-1}\times\mathbb{R}})$, the subspace of even finite (Radon) measures on ${\mathbb{S}^{d-1}\times\mathbb{R}}$. The arguments of the proof are based on duality. Indeed, observe that the adjoint of $\operatorname{T}_{\operatorname{ReLU}}$ takes the form $\operatorname{T}_{\operatorname{ReLU}}^{*}=\Delta\operatorname{\mathscr{R}}^{*}\operatorname{K}$ (since $\Delta$ and $\operatorname{K}$ are self-adjoint). Furthermore, from Proposition 4 combined with the fact that $\Delta:\mathcal{S}(\mathbb{R}^{d})\to\mathcal{S}(\mathbb{R}^{d})$ is continuous, we see that $\operatorname{T}_{\operatorname{ReLU}}^{*}:\mathcal{S}_{\operatorname{\mathscr{R}}}\to\mathcal{S}(\mathbb{R}^{d})$ is continuous. Therefore, by duality, $\operatorname{T}_{\operatorname{ReLU}}:\mathcal{S}^{\prime}(\mathbb{R}^{d})\to\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}$ is continuous. Since $r_{({\bm{w}},b)}\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$, we have that $\operatorname{T}_{\operatorname{ReLU}}r_{({\bm{w}},b)}$ is indeed well-defined. Finally, $\mathcal{M}^{\mathsf{e}}({\mathbb{S}^{d-1}\times\mathbb{R}})$ is continuously embedded in $\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}$ and so any finite measure in the range of $\operatorname{K}\operatorname{\mathscr{R}}$ can be concretely identified to have even symmetries (see Unser, 2023, for a detailed discussion). These symmetries are evidenced by the fact that the Radon transform of a “classical” function is necessarily even from the integral form in Eq. 13.
@@
    $\mathbf{P}(N_{\Pi}=n)=\frac{(\lambda\lvert\Pi\rvert)^{n}}{n!}e^{-\lambda\lvert\Pi\rvert},$ (21)
@@
$=\operatorname{ReLU}({\bm{u}}^{\mathsf{T}}{\bm{x}}-t)+{\bm{u}}_{0}^{\mathsf{T}}{\bm{x}}+t_{0},$ (26)
@@
where $\operatorname{T}_{\operatorname{ReLU}}^{\dagger*}$ is the limiting operator as $\varepsilon\to 0$ whose kernel is $k_{\bm{x}}\coloneqq\lim_{\varepsilon\to 0}k_{\bm{x}}^{\varepsilon}$. This random neural network is the unique CPwL solution to the SDE
@@ 7 Conclusion
We have investigated the statistical properties of random ReLU neural networks. We proved that these networks are well-defined non-Gaussian processes in the non-asymptotic regime. We showed that these processes are isotropic and wide-sense self-similar with Hurst exponent $3/2$. Remarkably, the autocovariances of these processes have simple closed-form expressions. Finally, we showed that, under suitable hypotheses, as the expected width tends to infinity, these processes can converge in law not only to Gaussian processes, but also to non-Gaussian processes depending on the law of the weights. These asymptotic results recover the classical observation that wide networks converge to Gaussian processes as well as prove that wide networks can converge to non-Gaussian processes under certain choices of initialization.
@@
This work was supported in part by the Swiss National Science Foundation under Grant 200020_219356 / 1 and in part by the European Research Council (ERC Project FunLearn) under Grant 101020573.
@@ Appendix A Proof of Theorem 9
is the null space of the Laplacian operator. It is well-known that $\mathcal{N}_{\Delta}$ is infinite-dimensional and that its members are necessary polynomials, the so-called harmonic polynomials. Therefore, the members of $\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ are actually equivalence classes of the form
@@
specified in Eq. 33 continuously maps $\mathcal{S}_{\Delta}(\mathbb{R}^{d})\to\mathcal{S}_{\operatorname{\mathscr{R}}}$ (cf., Unser, 2023, p. 21). Observe that, by Proposition 7, its extension by duality $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon}:\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}\to\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ coincides with $\operatorname{T}_{\operatorname{ReLU}}^{-}$. In particular, $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon}$ imposes the boundary conditions from Eq. 24 on the affine component of the harmonic polynomials in the equivalence classes in $\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$. Said differently, the range space $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon}\mathopen{}\mathclose{{}\left(\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}}\right)$ is the closed subspace of $\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ whose equivalence class members $[s]\in\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ additionally satisfy
@@
for any $\varphi\in\mathcal{S}(\mathbb{R}^{d})$, where $C=\lambda\,{\mathbf{E}}[\lvert V\rvert]<\infty$, where in the last line we used Eq. 74 with $p=1$. Since $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon*}:\mathcal{S}(\mathbb{R}^{d})\to L^{1}({\mathbb{S}^{d-1}\times\mathbb{R}})$ linearly and continuously, Proposition 3.1 of Fageot and Unser (2019) then guarantees that $\widehat{\mathbf{P}}_{s_{\operatorname{ReLU}}}$ is continuous, positive definite, and satisfies $\widehat{\mathbf{P}}_{s_{\operatorname{ReLU}}}(0)=1$. Therefore, the Bochner–Minlos theorem ensures that $\widehat{\mathbf{P}}_{s_{\operatorname{ReLU}}^{\varepsilon}}$ is the characteristic functional of the well-defined stochastic process $s_{\operatorname{ReLU}}^{\varepsilon}$.
By taking $\varepsilon\to 0$, we see that the random neural network $s_{\operatorname{ReLU}}\sim\mathcal{RP}(\lambda;\mathbf{P}_{V})$ is a measurable mapping $(\Omega,\mathcal{F},\mathbf{P})\to(\mathcal{S}^{\prime}(\mathbb{R}^{d}),\operatorname{\mathcal{B}}(\mathcal{S}^{\prime}(\mathbb{R}^{d}))$ whose characteristic functional is
@@
is continuous, remains to be true since $k_{\bm{x}}=\lim_{\varepsilon\to 0}k^{\varepsilon}_{\bm{x}}$ is compactly supported. Consequently, $s_{\operatorname{ReLU}}$ is the unique CPwL solution to SDE Eq. 36.

### v2

@@ 1 Introduction
A shallow (single-hidden-layer) neural network is a function of the form
@@ 2 Generalized Stochastic Processes
The mathematical framework used in this paper is based on the theory of generalized stochastic processes (Itô, 1954; Gelfand, 1955; Gelfand and Vilenkin, 1964; Itô, 1984) as opposed to the more common “time-series” approach to studying stochastic processes. In this section, we present the relevant background on generalized stochastic processes. We also refer the reader to the book of Unser and Tafti (2014) for further background. While this theory relies on some rather heavy concepts from functional analysis, it allows for elegant arguments to investigate the properties of the stochastic processes realized by the random neural networks in Eqs. 2 and 3.
@@ 4 Random ReLU Neural Networks as Stochastic Processes
where $\widetilde{{\bm{w}}}={\bm{w}}/\lVert{\bm{w}}\rVert_{2}$ and $\widetilde{b}=b/\lVert{\bm{w}}\rVert_{2}$. Therefore, the space of functions representable by shallow ReLU neural networks with input weights constrained to be unit norm is the same as the space of functions representable by shallow ReLU neural networks without constraints on the weights (Parhi and Nowak, 2023b; Shenouda et al., 2024). To that end, we focus on neurons of the form $\operatorname{ReLU}({\bm{w}}^{\mathsf{T}}{\bm{x}}-b)$ with $({\bm{w}},b)\in{\mathbb{S}^{d-1}\times\mathbb{R}}$.
@@
The equality in Eq. 18 is understood in $\mathcal{M}^{\mathsf{e}}({\mathbb{S}^{d-1}\times\mathbb{R}})$, the subspace of even finite (Radon) measures on ${\mathbb{S}^{d-1}\times\mathbb{R}}$. The arguments of the proof are based on duality. Indeed, observe that the adjoint of $\operatorname{T}_{\operatorname{ReLU}}$ takes the form $\operatorname{T}_{\operatorname{ReLU}}^{*}=\Delta\operatorname{\mathscr{R}}^{*}\operatorname{K}$ (since $\Delta$ and $\operatorname{K}$ are self-adjoint). Furthermore, from Proposition 4 combined with the fact that $\Delta:\mathcal{S}(\mathbb{R}^{d})\to\mathcal{S}(\mathbb{R}^{d})$ is continuous, we see that $\operatorname{T}_{\operatorname{ReLU}}^{*}:\mathcal{S}_{\operatorname{\mathscr{R}}}\to\mathcal{S}(\mathbb{R}^{d})$ is continuous. Therefore, by duality, $\operatorname{T}_{\operatorname{ReLU}}:\mathcal{S}^{\prime}(\mathbb{R}^{d})\to\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}$ is continuous. Since $r_{({\bm{w}},b)}\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$, we have that $\operatorname{T}_{\operatorname{ReLU}}r_{({\bm{w}},b)}$ is indeed well-defined. Finally, $\mathcal{M}^{\mathsf{e}}({\mathbb{S}^{d-1}\times\mathbb{R}})$ is continuously embedded in $\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}$ and so any finite measure in the range of $\operatorname{K}\operatorname{\mathscr{R}}$ can be concretely identified to have even symmetries (see Unser, 2023; Parhi and Unser, 2024, for a detailed discussion). These symmetries are evidenced by the fact that the Radon transform of a “classical” function is necessarily even from the integral form in Eq. 13.
@@
$=\operatorname{ReLU}({\bm{u}}^{\mathsf{T}}{\bm{x}}-t)+{{\bm{u}}_{0}^{\varepsilon}}^{\mathsf{T}}{\bm{x}}+t_{0}^{\varepsilon},$ (26)
@@
where $\operatorname{T}_{\operatorname{ReLU}}^{\dagger*}$ is the limiting operator as $\varepsilon\to 0$ whose kernel is $k_{\bm{x}}\coloneqq\lim_{\varepsilon\to 0}k_{\bm{x}}^{\varepsilon}$ (pointwise limit). This random neural network is the unique CPwL solution to the SDE
@@ 5 Properties of Random ReLU Neural Networks
    ${\mathbf{E}}[s_{\operatorname{ReLU}}({\bm{x}})]=\lambda{\mathbf{E}}[V]\int_{\mathbb{R}}\int_{\mathbb{S}^{d-1}}k_{{\bm{x}}}({\bm{u}},t)\,\mathrm{d}{\bm{u}}\,\mathrm{d}t,$ (39)
    where $k_{{\bm{x}}}$ is defined in Theorem 9.
@@ 7 Conclusion
We have investigated the statistical properties of random ReLU neural networks. We proved that these networks are well-defined non-Gaussian processes in the non-asymptotic regime. We showed that these processes are isotropic and wide-sense self-similar with Hurst exponent $3/2$. Remarkably, the autocovariances of these processes have simple closed-form expressions. Finally, we showed that, under suitable hypotheses, as the expected width tends to infinity, these processes can converge in law not only to Gaussian processes, but also to non-Gaussian processes depending on the law of the weights. These asymptotic results recover the classical observation that wide networks converge to Gaussian processes as well as prove that wide networks can converge to non-Gaussian processes. Although the presented investigation only considered shallow random ReLU neural networks, an important direction of future work would be to generalize our exact characterizations to deeper networks. To that end, the techniques developed by Zavatone-Veth and Pehlevan (2021) could provide a starting point for that investigation.
@@
The authors would like to thank the anonymous reviewers and the action editor for their careful reading of the manuscript. This work was supported in part by the Swiss National Science Foundation under Grant 200020_219356 / 1 and in part by the European Research Council (ERC Project FunLearn) under Grant 101020573.
@@ Appendix A Proof of Theorem 9
is the null space of the Laplacian operator. It is well-known that $\mathcal{N}_{\Delta}$ is infinite-dimensional and that its members are necessarily polynomials, the so-called harmonic polynomials. Therefore, the members of $\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ are actually equivalence classes of the form
@@
specified in Eq. 33 continuously maps $\mathcal{S}_{\Delta}(\mathbb{R}^{d})\to\mathcal{S}_{\operatorname{\mathscr{R}}}$ (cf., Parhi and Unser, 2025, Equation (A.3)). Observe that, by Proposition 7, its extension by duality $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon}:\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}\to\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ coincides with $\operatorname{T}_{\operatorname{ReLU}}^{-}$. In particular, $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon}$ imposes the boundary conditions from Eq. 24 on the affine component of the harmonic polynomials in the equivalence classes in $\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$. Said differently, the range space $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon}\mathopen{}\mathclose{{}\left(\mathcal{S}_{\operatorname{\mathscr{R}}}^{\prime}}\right)$ is the closed subspace of $\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ whose equivalence class members $[s]\in\mathcal{S}_{\Delta}^{\prime}(\mathbb{R}^{d})$ additionally satisfy
@@
for any $\varphi\in\mathcal{S}(\mathbb{R}^{d})$, where $C=\lambda\,{\mathbf{E}}[\lvert V\rvert]<\infty$, where in the last line we used Eq. 74 with $p=1$. Since $\operatorname{T}_{\operatorname{ReLU}}^{\dagger\varepsilon*}:\mathcal{S}(\mathbb{R}^{d})\to L^{1}({\mathbb{S}^{d-1}\times\mathbb{R}})$ linearly and continuously, Proposition 3.1 of Fageot and Unser (2019) then guarantees that $\widehat{\mathbf{P}}_{s_{\operatorname{ReLU}}^{\varepsilon}}$ is continuous, positive definite, and satisfies $\widehat{\mathbf{P}}_{s_{\operatorname{ReLU}}^{\varepsilon}}(0)=1$. Therefore, the Bochner–Minlos theorem ensures that $\widehat{\mathbf{P}}_{s_{\operatorname{ReLU}}^{\varepsilon}}$ is the characteristic functional of the well-defined stochastic process $s_{\operatorname{ReLU}}^{\varepsilon}$.
In the limiting scenario of $\varepsilon\to 0$, we see that the random neural network $s_{\operatorname{ReLU}}\sim\mathcal{RP}(\lambda;\mathbf{P}_{V})$ is a measurable mapping $(\Omega,\mathcal{F},\mathbf{P})\to(\mathcal{S}^{\prime}(\mathbb{R}^{d}),\operatorname{\mathcal{B}}(\mathcal{S}^{\prime}(\mathbb{R}^{d}))$ whose characteristic functional is
@@
is continuous, remains to be true since $k_{\bm{x}}=\lim_{\varepsilon\to 0}k^{\varepsilon}_{\bm{x}}$ (pointwise limt) is compactly supported. Consequently, $s_{\operatorname{ReLU}}$ is the unique CPwL solution to SDE Eq. 36.121212The mollifier argument is necessary in order to make sense of the boundary conditions Eq. 36 for elements of $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ that are not regular enough for the derivatives to exist.